OpenStudy (anonymous):

y varies directly with x. When y = 12, x = 16. Find x when y = 24. 18 32 64

mathslover (mathslover):

As y got doubled. So, x should also be doubled. Therefore, final value of x = double of original x = 2 * x = 2 * 16 = ?

OpenStudy (anonymous):

32

OpenStudy (anonymous):

For the next one, 18=12 * 1.5=1.5y, 1.5x = 1.5 * x = 1.5 * 16 = ?

OpenStudy (anonymous):

To solve for x when y has a certain value - let's call it a - find the ratio between 12 and a by dividing a by 12. Let's call the value of "a/12" b, multiply 16 by b to find the value of x when the value of y=a

OpenStudy (anonymous):

If that makes any sense